OpenStudy (anonymous):

Simply the expression. Assume that all variables are real numbers. 7squareroot 128(x-9)^7

OpenStudy (saifoo.khan):

\[laege \sqrt[7]{128(x-9)}^7\]Like that?

OpenStudy (anonymous):

yeah

OpenStudy (saifoo.khan):

\[\large \sqrt[7]{128(x-9)^7} \to [128(x-7)]^\frac77 \to [128(x-9)]^\frac{\cancel7}{\cancel{7}} \to 128(x-9)\]